OpenStudy (emogirlatemoooocow):

A box contains 4 oranges, 9 apples, and 3 pears. Liliana selects a piece of fruit from the box at random. What is the probability that she selects an orange or a pear?

Vocaloid (vocaloid):

hint: (number of oranges + pears)/(total number of fruit)

OpenStudy (emogirlatemoooocow):

A .\[\frac{ 1 }{ 20}\]

OpenStudy (emogirlatemoooocow):

B. \[\frac{ 3 }{ 64 }\]

OpenStudy (emogirlatemoooocow):

C. \[\frac{ 7 }{ 16 }\]

OpenStudy (emogirlatemoooocow):

D. \[\frac{ 9 }{ 16}\]

OpenStudy (myth1111):

How many oranges and pears are there together?

OpenStudy (emogirlatemoooocow):

12 ^

Vocaloid (vocaloid):

are you sure? read the problem again

OpenStudy (emogirlatemoooocow):

oops i added apples my bad

OpenStudy (emogirlatemoooocow):

7

OpenStudy (myth1111):

Right, now how many total fruit are there?

OpenStudy (emogirlatemoooocow):

add all the fruit together ? right :3

OpenStudy (myth1111):

Yeah

OpenStudy (emogirlatemoooocow):

okay the original is 4+9+3 is 16 7 + 9+3 is 19

OpenStudy (myth1111):

No, you just need the original

OpenStudy (myth1111):

So you have 7, and you have 16, so what's the answer?

OpenStudy (emogirlatemoooocow):

C. \[\frac{ 7 }{ 16 }\]

OpenStudy (myth1111):

Yup :)